# UGM2N: An Unsupervised and Generalizable Mesh Movement Network via M-Uniform Loss

### v1

@@ A.1 The proposed mesh movement method
A.2 Analysis of M-Uniform loss
@@
Moreover, from Eq. 9, we have $\mathcal{L}_{\mathrm{var}}(\mathcal{M})=\operatorname{Var}({L_{K}^{\prime}})$. When the samples in $L_{K}$ are repeated samples from $L_{K}^{\prime}$ (i.e., each sample is duplicated $n$ times), we have $\operatorname{Var}(L_{K})=\operatorname{Var}(L_{K}^{\prime})$. Assuming that each mesh element appears in approximately the same number of local patches—i.e., the sampling of $L_{K}^{\prime}$ in $L_{K}$ is nearly identical—we can adopt the approximation $\operatorname{Var}(L_{K})\approx\operatorname{Var}(L_{K}^{\prime})$. This assumption holds, for example, in high-quality triangular meshes generated by mesh generation software, where almost all mesh nodes have a degree of 6. Moreover, the number of nodes on the boundary is relatively small compared to the number of interior nodes. With $\operatorname{Var}\left(\mathbb{E}[L_{K}\mid I]\right)\geq 0$, we get such an inequality:
@@
Appendix B Dataset
B.1 Train data setups
@@
Figure 10: Flow flieds for trianing the mesh movement model.
B.2 Test data setups
@@
Moving cylinder and supersonic flow over the wegde. Both cases are benchmark cases from the Kratos official website [40]. The mesh element count for the moving cylinder case is 9,852, and the element count for the supersonic flow case is 27,115. Please refer to the official site for more details.
B.3 Monitor function setups
@@
Appendix C More model training details
@@
Appendix D Mesh adaptation results on Poisson’s equation and Burgers’ equation
Appendix E The mesh equidistribution condition on the adapted mesh
Solution of the Helmholtz equation
MA [37] 3.21e-7 3.89e-7 1.68e-7 1.30e-7 1.50e-7
M2N [9] 4.41e-7 5.90e-7 2.62e-7 4.67e-7 4.82e-7
UM2N [39] 5.41e-7 5.50e-7 2.89e-7 3.32e-7 3.73e-7
UGM2N 2.12e-7 2.68e-7 1.52e-7 1.10e-7 1.19e-7
Appendix F Evaluation metrics
@@
Appendix G Limitations and broader impacts
Limitations. 1) The model does not handle mesh nodes on the boundary; future work should consider how to achieve the movement of boundary mesh nodes. 2) The approximation in Eq. 12 relies on the assumption that "each mesh elment appears approximately the same number of times in local patches." However, this condition only holds when the degrees of mesh nodes are nearly identical. Therefore, the method still requires further testing on non-uniform meshes (e.g., where mesh node degrees vary significantly). 3) The model adopts a relatively simple and lightweight architecture. Future work could explore more complex model designs to achieve better mesh adaptation performance.

### v2

@@ B.1 The proposed mesh movement method
B.2 Analysis of M-Uniform loss
@@
Moreover, from Eq. 9, we have $\mathcal{L}_{\mathrm{var}}(\mathcal{M})=\operatorname{Var}({L_{K}^{\prime}})$. When the samples in $L_{K}$ are repeated samples from $L_{K}^{\prime}$ (i.e., each sample is duplicated $n$ times), we have $\operatorname{Var}(L_{K})=\operatorname{Var}(L_{K}^{\prime})$. Assuming that each mesh element appears in approximately the same number of local patches—i.e., the sampling of $L_{K}^{\prime}$ in $L_{K}$ is nearly identical—we can adopt the approximation $\operatorname{Var}(L_{K})\approx\operatorname{Var}(L_{K}^{\prime})$. This regularity assumption holds, for example, in high-quality triangular meshes generated by mesh generation software, where almost all mesh nodes have a degree of 6. Moreover, the number of nodes on the boundary is relatively small compared to the number of interior nodes. With $\operatorname{Var}\left(\mathbb{E}[L_{K}\mid I]\right)\geq 0$, we get such an inequality:
@@
Convergence and approximation analysis. The M-Uniform loss optimizes local mesh equidistribution by minimizing $\mathcal{L}_{\mathrm{var}}(\mathcal{P}_{i})$ (Eq. 7), similar to traditional mesh smoothing techniques that optimize geometric quality metrics (e.g., aspect ratio). In a serial implementation, sequential node updates improve $\mathcal{L}_{\mathrm{var}}(\mathcal{P}_{i})$ for a specific mesh node $i$ at each iteration, suggesting convergence. However, training errors limit exact minimization ($\mathcal{L}_{\mathrm{var}}(\mathcal{P}_{i})=0$), and serial updates are computationally inefficient. Our Jacobi-style parallel updates enhance efficiency but complicate analysis due to data races, update conflicts, and the nonlinear nature of the transformation represented by our model. The non-convex Hessian of the objective function further precludes standard optimization guarantees. Despite these theoretical challenges, we demonstrate through extensive experiments in Section 4 that UGM2N achieves robust convergence.
To further validate convergence, we track the nodal displacement norm $\delta_{k}=\|\mathbf{X}^{k+1}-\mathbf{X}^{k}\|$ over iterations, where $\mathbf{X}^{k}$ denotes the mesh node positions at iteration $k$. As shown in Fig. 11, $\delta_{k}$ exhibits consistent decay toward zero across five Helmholtz equation cases, confirming convergence. Although $\delta_{k}$ may stagnate or even increase for larger iterations ($k>10$) in some instances, our dynamic termination strategy (App. B.1) effectively mitigates this, ensuring stable mesh adaptation.
Figure 11: Nodal displacement norm $\delta_{k}=\|\mathbf{X}^{k+1}-\mathbf{X}^{k}\|$ across iterations for Helmholtz equations.
Appendix C Dataset
C.1 Train data setups
@@
Figure 12: Flow fields for trianing the mesh movement model.
C.2 Impact of training data volume
UGM2N attains robust generalization using merely 4 flow fields (yielding 10,440 node-patch samples) via its unsupervised learning paradigm. To quantify the effect of training data volume, we evaluated model performance across 4, 8, and 16 flow fields—corresponding to 10,440, 20,880, and 41,760 patches, respectively. As shown in Table 3, performance improves steadily with increasing data volume. Notably, the Poisson and Helmholtz cases exhibit substantial error reductions, improving from 5.21% to 7.43% and 9.94% to 14.21%, respectively. Overall, scaling the training data volume reliably boosts UGM2N’s generalization capability, with especially marked gains in complex PDEs, highlighting the critical role of larger datasets in future developments.
Table 3: Error reduction ( %) vs. training data volume (node patches)
Data Volume Poisson Helmholtz Burgers
10,440 5.21 9.94 30.07
20,880 6.56 13.68 29.36
41,760 7.43 14.21 30.56
C.3 Test data setups
@@
Moving cylinder and supersonic flow over the wedge. Both cases are benchmark cases from the Kratos official website [41]. The mesh element count for the moving cylinder case is 9,852, and the element count for the supersonic flow case is 27,115. Please refer to the official site for more details.
C.4 Monitor function setups
@@
Appendix D More model training details
@@
Appendix E Results
E.1 Evaluation metrics
@@
E.2 Mesh tangling evaluation on non-convex meshes
To further test whether our model avoids mesh tangling during the adaptation process on non-convex meshes, we adopted the same testing method as UM2N [39]. Specifically, we generated 100 meshes, each containing 8 nodes, with four located at the corners of a unit square and the remaining four randomly sampled within the unit square. The flow field was constructed using a mixture of Gaussian distributions. The results are shown in Table 5, demonstrating that both our model and UM2N consistently produced valid meshes. In contrast, the MA method struggles with non-convex cases and only generates valid meshes in limited scenarios. Fig. 13 shows some meshes after adaptation by our UGM2N model; our model generalizes well to unseen non-convex meshes without producing any mesh tangling.
Table 5: Results of the mesh tangling test on non-convex meshes
Metric MA UM2N UGM2N
[…]
Valid mesh 41 100 100
Figure 13: Mesh adaptation results of non-convex meshes (10 out of 100 cases).
E.3 Mesh adaptation results on Poisson’s equation and Burgers’ equation
E.4 Impact of boundary node treatment on model performance
Ensuring geometric consistency of mesh boundaries remains a key challenge in learned mesh movement. The core difficulty arises from inevitable prediction errors in ML models: even minor deviations in predicted boundary node positions can severely distort boundary layers—the thin near-wall regions where viscous, thermal, or electromagnetic gradients are physically dominant and numerically critical.
To further evaluate different boundary handling strategies, we conducted experiments enabling simple clip-to-boundary projection (i.e., projecting adapted boundary nodes back onto the domain boundary if they drift outside), with results summarized in Table 6. The results show that clip projection yields modest gains in internally dominated flows—where gradients or Hessian values are predominantly large in the interior—but delivers substantial improvements when large Hessian values are concentrated near the boundaries while interior values remain small (e.g., Poisson4, Poisson6, Poisson7).
These findings confirm that, while our fixed-boundary strategy ensures stability and prevents boundary layer collapse, simple post-hoc projection can safely enhance performance in scenarios where flow features vary sharply near boundaries. This highlights considerable potential for future improvements through learned boundary constraints or hybrid adaptation-projection schemes, especially in wall-bounded and multi-physics applications.
Table 6: Error reduction with and without boundary clip projection. Cases follow Table 1 in the main paper.
Burgers0 Burgers1 Helmholtz1 Helmholtz2 Helmholtz3 Helmholtz4 Helmholtz5
ER w.o. Clip (%) 32.22 30.19 14.11 13.15 15.03 14.09 16.98
ER with Clip (%) 35.27 30.18 16.68 16.17 15.94 15.57 17.34
Gain (%) $\uparrow$3.05 $\downarrow$0.01 $\uparrow$2.57 $\uparrow$3.02 $\uparrow$0.91 $\uparrow$1.48 $\uparrow$0.36
Poisson1 Poisson2 Poisson3 Poisson4 Poisson5 Poisson6 Poisson7
ER w.o. Clip (%) 14.56 9.00 12.46 1.70 9.07 4.90 2.56
ER with Clip (%) 17.49 14.82 13.08 15.98 11.18 17.85 15.69
Gain (%) $\uparrow$2.93 $\uparrow$5.82 $\uparrow$0.62 $\uparrow$14.28 $\uparrow$2.11 $\uparrow$12.95 $\uparrow$13.13
E.5 The mesh equidistribution condition on the adapted mesh
Solution of the Helmholtz equation
MA [38] 3.21e-7 3.89e-7 1.68e-7 1.30e-7 1.50e-7
M2N [9] 4.41e-7 5.90e-7 2.62e-7 4.67e-7 4.82e-7
UM2N [40] 5.41e-7 5.50e-7 2.89e-7 3.32e-7 3.73e-7
UGM2N (ours) 2.12e-7 2.68e-7 1.52e-7 1.10e-7 1.19e-7
E.6 Detailed efficiency analysis and scalability
A key consideration is the scalability of our model to large meshes. Theoretically, our GNN-based model imposes no restrictions on mesh scale, as it can process graphs of arbitrary size. However, practical GPU memory constraints limit single-GPU performance. On an RTX TITAN GPU (24GB), UGM2N can process up to  40,000 mesh elements, with inference requiring 345ms and interpolation 61ms, achieving a significant speedup over the MA method (25,510ms).
To further evaluate scalability, we measured inference, interpolation, and I/O times (data transfer between CPU and GPU) across varying mesh sizes (Table 8).It can be observed that I/O operations account for the majority of the time overhead. This is because, in each epoch, the mesh must be transferred from the GPU back to the CPU for Delaunay triangulation-based interpolation. If a GPU-accelerated Delaunay triangulation algorithm were available, both inference and interpolation could be performed entirely on the GPU—this represents one of the key directions for future optimization.
Furthermore, UGM2N’s absence of data dependencies between node patches enables efficient parallelization across $N$ GPUs, with processing time scaling as $O(1/N)$ (excluding communication overheads). This supports real-time mesh adaptation for large-scale simulations.
Table 8: Timing breakdown (ms) for UGM2N across mesh sizes. For each element count, we repeated the experiment ten times and report the mean and standard deviation.
Elements Inference Interpolation I/O
[…]
E.7 The mesh adaptation results on the irregular and anisotropic mesh
In App. B.2, we derive the validity of the proposed loss function by assuming mesh regularity $\operatorname{Var}(L_{K})\approx\operatorname{Var}(L_{K}^{\prime})$. This assumption holds for meshes generated by mature mesh generation software, as they often ensure a certain level of mesh quality (e.g., for a 2D test case with 60k mesh elements generated by mature mesh generation software, we observed that 92.96% of nodes had a degree of 6, while 2.96% had a degree of 5—meaning around 95% of nodes fell into these two categories). Furthermore, we tested the model’s applicability to irregular and anisotropic meshes—i.e., scenarios where the assumption does not hold.
Irregular mesh. We uniformly sampled mesh points within a unit rectangular domain and generated meshes using Delaunay triangulation, followed by mesh smoothing to optimize quality. As shown in Fig. 16, the resulting meshes feature numerous nodes with varying degrees, thus failing to satisfy the regularity assumption. We generated a total of 10 such samples, with each sample’s flow field solution defined as $\cos(2\pi(x-\mu_{x}))$, where $\mu_{x}$ is a random value uniformly drawn from $[0,1]$. The error reduction results for our model are presented in Table 9. These results demonstrate that, despite being trained on regular meshes, our model generalizes effectively to highly irregular meshes, achieving substantial error reduction in 8 out of 10 cases.
Table 9: Error reduction of UGM2N on irregular meshes
Case 1 Case 2 Case 3 Case 4 Case 5 Case 6 Case 7 Case 8 Case 9 Case 10
ER (%) -2.07 16.07 10.69 11.11 18.10 14.24 -2.10 8.46 9.84 7.16
Figure 16: Mesh adaptation results on irregular meshes (successful cases). It can be observed that mesh nodes move toward regions of high Hessian values in the flow field.
Anisotropic mesh. To evaluate our model’s performance on anisotropic meshes, we solve the Poisson equation with the exact solution $u(x,y)=(1-e^{\frac{-(x-0.5)^{2}}{0.01}})((x-0.5)^{2}-1)$ for $x,y]\in[0,1]\times[0,1]$. We first generate preliminary stretched mesh elements via anisotropic adaptation to produce the initial anisotropic mesh, as shown in Fig. 17. This mesh is then input to the mesh adaptation models, which further optimize the node distribution to align with the solution’s characteristics. In this solution, the Hessian values peak at approximately $x=0.3$, $0.5$, and $0.7$. As illustrated in Fig. 17, compared to the initial anisotropic mesh, both the UM2N and UGM2N models more accurately capture these three locations, while the MA method yields an invalid mesh. The absolute error reductions achieved after solving with the optimized anisotropic meshes are presented in Table 10. These results demonstrate that our model achieves the optimal error reduction, confirming its applicability to anisotropic meshes.
Figure 17: Mesh adaptation results on Anisotropic meshes. The UGM2N model successfully captures the peak locations of the Hessian values.
Table 10: Absolute error reduction on anisotropic mesh
MA M2N UM2N UGM2N
Absolute error reduction Fail Fail 0.02 0.09
E.8 Large-scale robustness validation on random geometries
To rigorously evaluate generalization under extreme geometric and flow variability, we conducted a large-scale experiment on 1,000 randomly generated mesh samples. Each sample features a random polygon with 3–6 edges and a synthetic flow field drawn from a Gaussian mixture model. We solved the Helmholtz equation using these flow fields as exact solutions and compared mesh adaptation performance across methods.
Results are presented in Table 11 and Fig. 18. Both MA and UM2N exhibit high variance and frequent negative error reduction, reflecting poor generalization across diverse geometries and flow structures. In contrast, UGM2N achieves a Positive ER Ratio of 0.807 (807 successful cases out of 1,000) with a stable mean ER of 13.99% $\pm$ 21.05%, demonstrating superior robustness. M2N was excluded from comparison due to its inability to generalize to arbitrary geometries and flow fields without case-specific training.
Table 11: Robustness comparison on 1,000 random polygonal domains with Gaussian mixture flow fields (Helmholtz equation). Positive ER Ratio = fraction of cases with ER > 0.
Method ER (mean% $\pm$ std%) Positive ER Ratio
MA -227.76 $\pm$ 468.82 0.110
UM2N -50.72 $\pm$ 68.36 0.245
UGM2N (Ours) 13.99 $\pm$ 21.05 0.807
Figure 18: Mesh adaptation results of UGM2N on different flow fields across different meshes (30 cases out of 1000). Whether it is changes in the mesh or changes in the flow field, UGM2N demonstrates excellent robustness.
E.9 More results of qualitative experiments
For the supersonic flow and moving cylinder cases, MA method exhibited divergence during the solution of the monitor function equation in both cases, ultimately failing to produce valid adapted meshes. As shown in Fig. 19, the M2N method displayed inconsistent behavior, particularly in the moving cylinder case, where it generated well-adapted meshes at certain time steps but yielded highly distorted elements at others. For the supersonic flow, M2N successfully captured the shock position; however, the resulting mesh exhibited lower nodal density at the shock compared to UGM2N. Similarly, UM2N identified the shock location in the supersonic case but introduced undesirable element distortions in irrelevant regions away from the shock. In the moving cylinder scenario, while UM2N effectively fitted the flow field distribution, its excessive adjustments led to prominent mesh distortions.
Figure 19: Mesh adaptation results of the M2N and UM2N models for the supersonic flow and moving cylinder cases. Red boxes mark the locations of mesh distortions.
Appendix F Limitations and broader impacts
Limitations. 1) The model currently fixes boundary nodes during adaptation to ensure geometric consistency and prevent boundary layer distortions; while post-hoc clip projection to the boundary yields modest improvements in some cases, future work could integrate learned boundary constraints for more flexible handling. 2) The approximation in Eq. 13 assumes mesh regularity (i.e., similar node degrees), but empirical tests on irregular meshes (App. E.7) demonstrate robust performance in 80% of cases despite violations. Further validation on highly non-uniform or anisotropic meshes with extreme degree variations remains warranted. 3) The model adopts a relatively simple and lightweight architecture. Future work could explore more complex model designs to achieve better mesh adaptation performance. 4) By adopting a node-patch-based processing strategy, our method achieves scale invariance and translation invariance—that is, adaptation results remain unchanged under uniform scaling or translation of the mesh. However, we observe that all AI-based mesh adaptation methods currently lack rotation invariance: rotating the coordinate system of the mesh leads to inconsistent results. This remains an important open challenge for future work.